A solution is prepared by dissolving 17.75 g sulfuric acid, H2SO4, in enough water make 100.0 mL. What is the molarity? A) 1.940 M H2SO$ B) 0.1810 M H2SO4 C) 1.810 M H2SO4 D) 0.1775 M H2SO4
Answer #1

H2SO4 molar mass=98
moles of H2SO4 = 17.75/98 = 0.1811

molarity of H2SO4=moles /volume = 0.1811/0.1L = 1.810 M

hence answer is C) 1.810M H2SO4